### v2

@@ Abstract
Data and Code Availability
We use the following public datasets: MIMIC-CXR-JPG (johnson2019mimic), MS-CXR (boecking2022making), and MS-CXR-T (bannur2023learning), which are accessible under their respective data use agreements. We plan to release the code after obtaining organizational approval.
Institutional Review Board (IRB)
This work does not require IRB approval.
@@ 1 Introduction
Chest X-ray (CXR) imaging is a cornerstone of pulmonary screening, diagnosis, and follow-up, accounting for the largest share of diagnostic radiology examinations performed worldwide each year (cid2024development). Yet systematic assessment of thoracic structures remains labor-intensive, imposing a substantial time burden on radiologists (fallahpourmedrax). The gradual introduction of AI into clinical practice shows promise for alleviating this workload (zhang2024generalist; tanno2025collaboration). However, prevailing systems fall short on complex multimodal reasoning, such as integrating findings across disparate image regions, which is central to radiologists’ practice; most adhere to end-to-end designs in which the visual encoder executes a single, front-end pass and subsequent reasoning proceeds purely in text (wang2025multimodal). This encode-once, text-only paradigm decouples the reasoning trajectory from evolving visual evidence, leading to failures on tasks that require iterative re-inspection, precise measurements, and cross-comparisons (liu2025more).
@@
A promising path for clinical reasoning is to augment large language models, including multimodal variants, with external tools (lu2025octotools). By delegating perceptual and classification subtasks such as organ or region segmentation and disease classification to validated modules, the language model can focus on planning and synthesis. Several agentic frameworks already explore this idea, ranging from training small models for limited tool use (li2024mmedagent; nath2025vila) to pipeline systems that invoke general-purpose models for more flexible operations (jiang2025medagentbench; schmidgall2024agentclinic), although multiple agent coordination and/or debate introduces considerable computational overhead. In CXR interpretation, RadFabric (chen2025radfabric) integrates diagnostic agents with a separate reasoning agent, and MedRAX (fallahpourmedrax) expands task coverage by incorporating additional task specific models. Despite gains over single-model baselines, integration and reasoning steps are often opaque and not aligned with clinical workflow, which undermines trust and creates safety risks, and they still lack explicit verification and conflict resolution.
To address these limitations, we introduce RadAgents, a framework designed for complex multimodal reasoning in CXR, which has three primary technical contributions:
  • We introduce clinically aware interleaved reasoning, a multimodal loop that combines clinical priors and study metadata with targeted image operations (segment, measure, localize, compare), so each hypothesis triggers tool use and returns inspectable evidence, mirroring how radiologists proceed from observation to measurement, comparison with prior studies, and conclusion (Fig. 6 and Appendix E).
  • We propose a training-free multi-agent design, where five sub-agents operate in clean, task-scoped context compartments, coordinated by an Orchestrator and reconciled by a Synthesizer. This preserves visual grounding, supports parallel execution, and composes measurement, localization, characterization, comparison, and diagnosis (Figure 1).
  • We add a lightweight context verifier and visual retrieval-augmented generation to detect, surface, and resolve inconsistencies before reporting, yielding more clinically aligned outputs.
  • RadAgents achieves state-of-the-art performance, outperforming strong baselines by 10.2% on MS-CXR, 29.6% on MIMIC-CXR, and 21.5% on MS-CXR-T. Multimodal retrieval further contributes an average 8.0% boost over the no-retrieval variant, mitigating context conflicts and improving reliability.
@@
2 Methods
RadAgents is a multi agent system with seven specialized agents (Figure 1). Five implement the clinical ABCDE review scheme (hodler2019diseases): Airway, Breathing, Circulation, Diaphragm, and Everything else. In addition, an Orchestrator agent analyzes each query and routes tasks to the appropriate specialists with the required patient context (for example, imaging view and prior studies), and a Synthesizer agent integrates their outputs, resolves conflicts, and produces the final output.
This design confines context to task specific compartments, reducing the information each agent must process and simplifying context compression by having each sub-agent produce an initial summary for downstream synthesis. It also allows parallel execution, lowering latency for long reasoning. For common and clinically significant CXR findings such as cardiomegaly and pleural effusion, we curate radiologist-like workflows, the predefined templates within RadAgents, to guide tool selection and clinically grounded reasoning (see the demonstration example in Appendix A). For out-of-template queries, the system invokes workflow-free reasoning, preserving flexibility. The design is extensible: new templates (e.g., reasoning or tool-chains) can be added, and some can generalize to tasks of similar scope or category.
2.1 Task-aware Sub-Agents
Each sub-agent, also called the ABCDE agent, has a defined purpose and domain of expertise. Each is governed by a custom system prompt derived from clinical guidelines and maintains its own context window (See details ). The main scope and objectives of them are:
@@ 2.1 Task-aware Sub-Agents
Breathing agent: Survey the lungs and pleura for parenchymal and pleural pathology; for example, detect opacities (alveolar, interstitial, nodular), infiltrates, and consolidation patterns.
@@
Everything Else agent: Identify other chest wall, soft tissue, and foreign body findings.
2.2 Radiologist-like Workflows
Solving different radiological tasks requires distinct reasoning modes and corresponding evidence. Figure 6 shows that multimodal reasoning, particularly interleaved quantitative measurements, strengthens reasoning for enlarged heart detection, whereas a simple grounding guided approach fails. Motivated by this, RadAgents curates clinically driven subtasks and workflow templates within the subagents’ system prompts, organized into five modes: (M1) measurement, (M2) localization, (M3) characterization, (M4) relational and comparative reasoning, and (M5) diagnosis. Complex clinical reasoning composes these modes and triggers the appropriate tool calls within the agentic system; for example, judging progression of effusion volume uses M1 and M4. A detailed description of workflows can be found in Appendix E.
@@ 2.3 Global Controller Module
Query analysis. Given a query, the Orchestrator drafts a high level plan, selects the relevant subagents, and chooses the reasoning mode: ReAct when no workflow is specified (yao2023react), or Plan-and-Execute (P&E) (wang2023lim) when a workflow template is available. This keeps the system language driven and adaptable across queries.
Tools. We employ a suite of models as tools for distinct CXR tasks: CheXagent (chen2024chexagent) for VQA, MAIRA-2 (bannur2024maira) for grounding, the CheXpert Plus report generator (chambon2024chexpert), and classification and organ segmentation models from TorchXRayVision (cohen2022torchxrayvision). In addition, we include unique programming tools that return zoomed-in quarter patches or serve for measurement and calculation purposes.
Context verifier. No tool is perfect, as their capabilities are constrained by model size and training data. When uncertainty arises, we trigger a verification step in which an advanced multimodal LLM serves as a judge (chen2024mllm), filtering out incorrect outputs such as erroneous masks.
Retrieval-augmented conflict resolution. Tool outputs can conflict. On the Synthesizer side, we apply Visual Retrieval-Augmented Generation (V-RAG) (chu2025reducing): the agent retrieves clinically similar chest radiographs (based on image embeddings from Rad-DINO (perez2025exploring)) and accompanying context like patient notes and uses these exemplars to adjudicate discrepancies among tools (Figure 2 and Appendix F). This mirrors routine radiologic practice, in which clinicians consult similar cases and content to calibrate interpretation.
@@ 3.1 Experimental Setup
To demonstrate the generality of RadAgents, we evaluate it on three tasks with increasing reasoning complexity: VQA for existence and attributes (E&A), VQA for comparison and progression (C&P), and report generation. The data statistic and details are shown in Table 2.
@@
Baselines. We instantiate all agents with GPT-4o. For comparison, we include (1) GPT-4o (hurst2024gpt), (2) GPT-4o with ReAct, where tools are available via function calling but without explicit workflow steering, (3) GPT-4o with monolithic Workflow where a MLLM executing ABCDE analysis end-to-end with the same tools, retrieval, and budges as RadAgents, and (4) two medical specialist models, CheXagent and MedGemma (sellergren2025medgemma). Unless otherwise noted, the number of retrieved exemplars for V-RAG is set to $k=3$ (see Appendix F for an ablation on $k$). We report results for two variants of RadAgents, with and without V-RAG.
Metrics. For VQA (E&A) and report generation, we use standard CXR text metrics (explanation in Appendix B): RadGraph F1 (jain1radgraph), CheXbert macro F1 across 14 labels (smit2020combining), RaTE (zhao2024ratescore), and GREEN (ostmeier2024green). Because the outputs are sentences, these metrics capture clinical correctness and consistency. For VQA (C&P), where the output is one of three choices, we report accuracy.
Table 1: Evaluation on VQA (E&A) and report generation with baselines and RadAgents. The four metrics are commensurate (each normalized to the range $[0,1]$).
Method CheXbert-macro-F1(14) RadGraph-F1 RaTE GREEN Avg.

### v3

@@ 1 Introduction
Chest X-ray (CXR) imaging is a cornerstone of pulmonary screening, diagnosis, and follow-up, accounting for the largest share of diagnostic radiology examinations performed worldwide each year (cid2024development). Yet systematic assessment of thoracic structures remains labor-intensive, imposing a substantial time burden on radiologists (fallahpourmedrax). The gradual introduction of AI into clinical practice shows promise for alleviating this workload (zhang2024generalist; tanno2025collaboration). However, prevailing systems fall short on complex multimodal reasoning, such as integrating findings across disparate image regions, views, and time points, which is central to radiologists’ practice. Most methods adhere to end-to-end designs in which the visual encoder performs a single front-end pass and subsequent reasoning proceeds purely in text (wang2025multimodal). This encode-once, text-only paradigm decouples the reasoning trajectory from evolving visual evidence, leading to failures on tasks that require iterative re-inspection, precise measurements, and cross-comparisons (liu2025more) as shown in Figure 1.
@@
A promising direction is to augment large language models, including multimodal variants, with external tools (lu2025octotools). By delegating perceptual and classification subtasks such as organ or region segmentation and disease classification to validated modules, the language model can focus on planning and synthesis. Several agentic frameworks have explored this idea, ranging from training small models for limited tool use (li2024mmedagent; nath2025vila) to pipeline systems that invoke general-purpose models for more flexible operations (jiang2025medagentbench; schmidgall2024agentclinic). In CXR interpretation, RadFabric (chen2025radfabric) integrates diagnostic agents with a separate reasoning agent, and MedRAX (fallahpourmedrax) expands task coverage by incorporating additional task-specific models. Despite improvements over single-model baselines, these systems often remain opaque and weakly aligned with clinical workflow: integration and reasoning steps are not explicitly traceable, visual and textual evidence are loosely coupled, and inconsistencies across tools are not systematically detected or resolved, which undermines trust and creates safety risks.
By contrast, radiologists reason through structured, radiology-specific workflows. For CXR, training and guidelines emphasize systematic review schemes, explicit quantitative assessments (e.g., cardiothoracic ratio, carinal angle), and comparison across time and views (hodler2019diseases). This process is inherently interleaved: clinicians move back and forth between image inspection, measurements, and textual synthesis, refining hypotheses as new evidence is obtained. Crucially, the reasoning is explicit and traceable, enabling peer review, error analysis, and integration with broader clinical context. Current multimodal LLM systems capture this paradigm only implicitly: reasoning is buried inside the model, tool calls are ad hoc, and there is limited support for auditing how a conclusion was reached or which intermediate steps failed (lee2025cxreasonbench).
To bridge this gap, we present RadAgents, a multi-agent framework for complex multimodal reasoning in CXR that encodes radiologist-like workflows into a modular, auditable pipeline. RadAgents decomposes interpretation into seven specialized agents: five subagents that implement core radiologic review modes, an Orchestrator that analyzes queries and dispatches tasks, and a Synthesizer that aggregates outputs, performs contextual verification, and resolves cross-tool conflicts. Each subagent follows predefined, radiologist-style workflows when applicable (e.g., for cardiomegaly or pleural effusion), while out-of-template queries fall back to workflow-free ReAct-style reasoning. Throughout, RadAgents maintains explicit logs of intermediate artifacts (segmentations, measurements, retrieved exemplars, and rationales), yielding step-level traceability instead of a single opaque explanation.
From a deployment standpoint, we instantiate RadAgents with open-source, lightweight vision–language models (Qwen3-VL-Instruct 4B/8B/30B (Qwen3-VL)) as the core engines of each agent. We show that, when paired with structured workflows, tool integration, and conflict resolution, an 8B open model can match or surpass GPT-4o and specialist CXR models such as CheXagent (chen2024chexagent) on diverse benchmarks. This suggests that trustworthy, workflow-aligned CXR reasoning does not strictly require very large proprietary models and can be realized with more accessible, on-premise-friendly architectures. Our contributions are threefold:
  • We formalize radiologist-like multimodal workflows for CXR interpretation and encode them in a multi-agent system that interleaves visual evidence, measurements, and textual reasoning. RadAgents supports both workflow-guided and workflow-free modes, enabling coverage of common guideline-style tasks as well as open-ended queries.
  • We design a traceable, tool-augmented agentic architecture that combines an Orchestrator, subagents, and a Synthesizer with retrieval-augmented conflict resolution and short-term memory. This yields explicit step-by-step trajectories and principled handling of cross-tool inconsistencies, improving alignment with clinical practice.
  • We conduct an extensive study on across three challenging multimodal medical-reasoning datasets. RadAgents consistently outperforms competitive baselines by a substantial margin, and ablations show the importance of radiologist-like workflows, visual retrieval, and targeted scaling of the Synthesizer.
2 Methodology
@@
RadAgents is a multi agent system with seven specialized agents (Figure 2). Five implement the clinical ABCDE review scheme (hodler2019diseases): Airway, Breathing, Circulation, Diaphragm, and Everything else. In addition, an Orchestrator agent analyzes each query and routes tasks to the appropriate specialists with the required patient context (for example, imaging view and prior studies), and a Synthesizer agent integrates their outputs, resolves conflicts, and produces the final output. This design confines context to task specific compartments, reducing the information each agent must process and simplifying context compression by having each sub-agent produce an initial summary for downstream synthesis. It also allows parallel execution, lowering latency for long reasoning.
2.1 Tool Set
To support radiologist-like reasoning, RadAgents integrates a comprehensive tool set that includes state-of-the-art machine learning models for specific tasks, general-purpose data processing utilities, and Python modules for measurement and calculation based on intermediate results, and for data utility.
  • ROI Segmentation. Region-of-interest (ROI; e.g., anatomical structures and lesions) segmentation plays a central role in medical reasoning, as it provides interpretable visual evidence and often constitutes the first step in diagnostic workflows. For example, cardiothoracic ratio (CTR) calculation requires measuring both thoracic width and cardiac width from segmentation masks.
    Tool list: (a) CXAS, an anatomy segmentation model (seibold2023accurate), which can segment up to 157 anatomical structures relevant to chest radiography; (b) BiomedParser (zhao2024biomedparse), a text-driven medical image parsing model that covers 82 major biomedical object ontologies, such as viral pneumonia.
  • Phrase Grounding. Unlike object-level segmentation, phrase grounding aims to localize a finding described by free-text (e.g., “right lower lobe opacity”) via a bounding box, providing finding-level evidence for generated outputs.
    Tool list: MAIRA-2 (bannur2024maira), selected because it is trained on diverse (public and private) grounded datasets.
  • Measurement and Calculation. Radiologists routinely assess the size, shape, and geometric relationships of ROIs to refine diagnoses, such as measuring the carinal angle or estimating pleural effusion volume for severity assessment.
    Tool list: We implement a suite of reusable Python scripts that perform geometric measurements and numeric calculations given either image inputs (e.g., segmentation masks, keypoints) or structured text. Further details are provided in Appendix B.
  • Visual Question Answering (VQA). Medical VQA models enable the agentic system to handle flexible free-form queries, especially when it is unnecessary or overly costly to execute full visual reasoning pipelines.
    Tool list: We adopt MedGemma (sellergren2025medgemma), which combines strong instruction-following ability with medical knowledge. As a specialist CXR model, CheXagent is added as a complementary tool.
  • Report Generation. Report generation models serve as references to initialize or supplement the final radiology report, particularly for routine findings.
    Tool list: We use the CheXpert Plus report generator (chambon2024chexpert). MAIRA-2 is reused here to provide grounded visual evidence that can be integrated into the generated report.
  • Pathology Classification. For certain pathology-specific pixel patterns, it is difficult to explicitly quantify the features needed for reasoning, and classification models become particularly valuable.
    Tool list: (a) A DenseNet-121 model from TorchXRayVision (cohen2022torchxrayvision), trained on four large-scale CXR datasets; (b) the VQA models, which can also be used in a classification mode by constraining their outputs.
  • Data Processing. General data-processing utilities include a DICOM loader (with metadata parsing for fine-grained measurement), visualization tools, and basic preprocessing operations such as contrast adjustment and resizing, which standardize inputs for downstream tools.
2.2 Task-aware Subagents
Each subagent, also called the ABCDE agent, has a defined purpose and domain of expertise. Each is governed by a custom system prompt and maintains its own context window. The main scope and objectives of them are:
@@ 2.2 Task-aware Subagents
Breathing agent: Survey the lungs and pleura for parenchymal and pleural pathology; for example, detect opacities (atelectasis, inltrate) and nodule.
@@
Everything-else agent: Identify chest wall (ribs and fractures), soft tissue, and foreign materials like medical devices.
Each subagent is equipped with an LLM and an individual skill set encoded in its system prompt. The skill set is a collection of reusable skill units, each defined as a reference tool chain together with decision thresholds for a specific clinical purpose. For example, computing the carinal angle requires first obtaining a segmentation mask of the tracheal bifurcation, then applying a geometric algorithm to identify the carina and main bronchi, and finally comparing the resulting angle against the normal range of 40–80 degrees.
The operational plan, i.e., which skills to invoke and in what order, is determined by the high-level intent distributed by the Orchestrator. Conditioned on this intent, each subagent performs step-by-step reasoning and tool use following the ReAct (“Reason + Act”) paradigm (yao2023react); if a step fails or produces inconsistent evidence, the subagent triggers local re-planning to repair the failure. This behavior is agentic rather than a fixed automation script, making the system more robust to uncertainty while still leveraging established clinical practice patterns. Details of the skill sets for each subagent are provided in Appendix B and A.
@@ 2.3 Global Controller Module
Query analysis. Given a query, the Orchestrator first analyzes the intent, extracts key clinical entities and objectives, and then drafts a high-level plan. It activates only the associated subagents (all other agents remain idle, incurring no additional computation cost) and sends them clear, goal-oriented instructions that specify the required clinical indicators without prescribing which tools to use. For example, for the query “Is there lung opacity?”, the Orchestrator generates a plan derived from our predefined Workflow, such as: “Goal: (1) determine the existence of lung opacity; (2) if present, determine the type; (3) determine the location; and (4) verify the answer.”
This decoupling between the Orchestrator and the tool set greatly improves maintainability: otherwise, adding a new tool or updating an existing one would require rewriting the workflow. Instead, we introduce a skill layer as an intermediate abstraction and let each subagent, guided by language understanding, dynamically compose tools to accomplish a given skill. In this way, workflows, skills, and tools can be maintained and evolved independently. If a task is dispatched incorrectly, the receiving subagent raises a SkillMismatchError to request re-dispatch.
@@
Retrieval-augmented conflict resolution. No tool is perfect, as its capabilities are bounded by model size and training data, and different tools may produce conflicting outputs. On the Synthesizer side, we therefore apply Visual Retrieval-Augmented Generation (V-RAG) (chu2025reducing): the Synthesizer retrieves clinically similar chest radiographs (using image embeddings from Rad-DINO (perez2025exploring)) together with associated context such as patient notes, and leverages these exemplars to adjudicate discrepancies among tools (Figure 3 and Appendix D). This design mirrors routine radiologic practice, where clinicians consult prior cases and reference material to calibrate their interpretations.
Short-term memory. RadAgents maintains a shared short-term memory that caches patient-specific context, including demographics, clinical indications, acquisition information (e.g., AP vs. PA views), and metadata when DICOM images are provided. In addition, the short-term memory stores tool outputs, which are accessible to all agents. If an agent needs to call a tool whose result is already cached, it directly reads the cached output instead of re-invoking the tool. This mechanism prevents redundant computation, reduces latency, and improves consistency in multi-step analyses that repeatedly reference the same intermediate results.
2.4 Implementation
RadAgents is implemented on top of the LangGraph agentic framework, and is conceptually inspired by MedRAX (fallahpourmedrax). The core engine of each agent can in principle be any LLM; in this work we focus on open-source, lightweight models, specifically Qwen3-VL-Instruct series, which exhibit strong instruction-following and reasoning capabilities and supports long-context input that is crucial for our agentic system. Because we explicitly encode clear, radiologist-style workflows into the agents, even the smaller models can reliably follow the prescribed procedures in most cases. To ensure transparency and enable debugging and analysis, we log the full execution trajectory and all intermediate outputs for each run.
For workflow extraction, we adopt hybrid retrieval over the textual workflow descriptions, combining BM25 with dense similarity based on the Snowflake-arctic-embed-m-v1.5 embeddings. Skills and tools are exposed as structured JSON APIs, and agents issue calls by constructing precise JSON objects that specify the target tool together with all required arguments (e.g., image paths and text prompts).
@@ 3.1 Experimental Setup
Datasets. To demonstrate the generality of RadAgents, we evaluate it on three benchmark datasets that closely mirror complex clinical workflows, offer sufficient task diversity, and are relatively easy to evaluate for correctness: (1) ChestAgentBench (fallahpourmedrax) includes 2,500 questions derived from expert-validated clinical cases, covering seven core competencies and associated reasoning skills essential for CXR interpretation. (2) For the CheXbench (chen2024chexagent) subset, following MedRAX, we focus on visual question answering (115 cases from Rad-Restruct (pellegrini2023rad) and 123 cases from SLAKE (liu2021slake)) and 380 fine-grained multimodal reasoning questions from OpenI  [^1] https://openi.nlm.nih.gov/. (3) We further evaluate on the preprocessed multi-view and longitudinal MIMIC-CXR test set (2,231 cases) used in EditGRPO (zhang2025editgrpo).
@@
Baselines. Unless otherwise specified, we instantiate all agents with Qwen3-VL-Instruct-8B. Additional results using GPT-4o as the agent core, with the same tool set as MedRAX, are provided in Appendix C. For comparison, we include: (1) specialist models CheXagent and GPT-4o; (2) Qwen3-VL in a single-agent setting using ReAct (following MedRAX) without workflow steering; (3) Qwen3-VL (single agent) using ReAct guided by our workflow templates (including the full skill set). Unless otherwise noted, the number of retrieved exemplars for V-RAG is set to $k=3$ (see Figure 7 and Appendix D for ablation study). We report results for two variants of RadAgents, with and without V-RAG.
Metrics. For ChestAgentBench and CheXbench, which consist of closed-ended questions, we report accuracy. For report generation, we use GREEN score (ostmeier2024green), which follows an LLM-as-a-judge paradigm and has been shown to align well with human judgments. Because the outputs are free-text sentences, this metric captures both clinical correctness and consistency.
3.2 Main Results
ChestAgentBench. Figure 4 compares the performance of different systems on the seven categories in ChestAgentBench. RadAgents achieves the best accuracy in every category, yielding an overall score of 73.6%, substantially higher than CheXagent (39.5%), GPT-4o (56.4%), Qwen3-VL w/ ReAct (61.3%), Qwen3-VL w/ Workflow (63.5%), and RadAgents without V-RAG (66.9%). The gains are consistent across tasks, with RadAgents outperforming the strongest non-agent baseline (Qwen3-VL w/ Workflow) by 771010 points on most categories. The largest margins are observed on diagnosis (73.9% vs. 64.3%) and characterization (71.0% vs. 61.5%), which require synthesizing subtle imaging findings and clinical priors. Comparing RadAgents with and without V-RAG also reveals a clear benefit from visual retrieval: V-RAG contributes around 6–7 absolute points overall, suggesting that access to external image evidence is particularly helpful for fine-grained and high-level reasoning questions.
Figure 4: Performance on ChestAgentBench across different categories of questions.
CheXbench. Table 1 reports results on CheXbench, which includes two VQA benchmarks (Rad-Restruct and SLAKE) and the OpenI image-text reasoning task. RadAgents again attains the highest overall accuracy (74.6%), improving upon both domain-specific CheXagent (64.7%) and the general-purpose GPT-4o (63.5%). On visual QA, RadAgents reaches 76.5% on Rad-Restruct and 89.4% on SLAKE, indicating strong capability in localized and fine-grained visual understanding. Qwen3-VL w/ Workflow and RadAgents w/o V-RAG are competitive, but RadAgents still provides a consistent 2–4 point advantage across all three sub-tasks. The OpenI reasoning task is more challenging for all models, yet RadAgents achieves 69.2% accuracy, outperforming RadAgents w/o V-RAG (66.3%) and other baselines. These results highlight that the proposed agentic workflow, together with visual retrieval, not only enhances structured VQA but also benefits more global image–text reasoning.
Table 1: Accuracy (%) comparison on CheXbench.
Model VQA OpenI Reasoning Overall
Rad-Restruct SLAKE
CheXagent 57.1 78.1 59.0 64.7
GPT-4o 53.9 85.4 51.1 63.5
Qwen3-VL w/ ReAct 70.4 86.2 61.3 68.0
Qwen3-VL w/ Workflow 72.2 87.8 65.3 71.0
RadAgents wo/ V-RAG 71.3 87.8 66.3 71.5
RadAgents 76.5 89.4 69.2 74.6
MIMIC-CXR Report Generation. Figure 5 reports GREEN scores on the MIMIC-CXR test set under the multi-view and longitudinal setting, which better reflects real clinical workflows. RadAgents attains the highest GREEN score of 51.4, outperforming RadAgents w/o V-RAG (46.1), GPT-4o w/ ReAct (42.3), GPT-4o w/ Workflow (41.7), plain GPT-4o (34.2), and CheXagent (23.6). The relatively marginal improvement of the workflow-based variants over plain GPT-4o suggests that simply feeding long multi-study contexts to a single agent is insufficient, as the model can become “lost in the middle” (liu2024lost) and under-utilize information dispersed across the sequence. The sizeable gap between RadAgents and its ablated variant highlights that visual retrieval and our specialized multi-agent design are particularly beneficial for generating temporally consistent reports conditioned on multiple studies.
Figure 5: Multi-view and longitudinal performance on MIMIC-CXR test set.
3.3 Ablation Study
We investigate how the capability of the core LLM (i.e., model scale/parameters) influences RadAgents by varying the backbone Qwen3-VL-Instruct model from 4B to 8B and 30B. Since RadAgents contains multiple roles, in each ablation we upgrade only a single component, either the Orchestrator or the Synthesizer, while keeping all other agents at the default 8B model [^2] Running all five agents with 30B models in parallel would be prohibitively expensive.. For example, when evaluating the Synthesizer, the Orchestrator and all subagents use the 8B model, whereas the Synthesizer uses the 30B model. This setup allows us to separately assess (a) whether a stronger Orchestrator can better orchestrate subagents (i.e., activate the correct specialists and deliver precise workflows), and (b) whether a stronger Synthesizer can more effectively resolve conflicts between tools and subagents.
The evaluation set comprises 100 representative cases: (a) 50 VQA instances randomly sampled from the MS-CXR test set, querying the existence and attributes of abnormalities (e.g., size and severity), and (b) 50 report-generation cases from the MIMIC-CXR test set, covering medium to high complexity. Details of the underlying datasets are provided in Appendix C.
Figure 7 shows that the dispatch success rate, defined as the proportion of cases where the correct subagents are activated and no “request re-dispatch” error is raised, increases from 87.0% to 93.0% as the Orchestrator model size grows from 4B to 30B. This indicates that stronger language understanding helps distribute tasks more reliably, though the gain is relatively modest because our hybrid search design already resolves most dispatch ambiguities. In contrast, the conflict-resolution rate of the Synthesizer (measured over the 38 cases with inter-tool disagreements, 9 from VQA and the remainder from report generation) improves substantially, rising from 26.3% (4B) to 44.7% (8B) and 60.5% (30B). These results suggest that conflict resolution is considerably more sensitive to model capacity than task dispatch, and that investing capacity in the Synthesizer is crucial for reliably reconciling heterogeneous tool and agent outputs.
Figure 6: figure
Effect of model scale on Orchestrator dispatch success and Synthesizer conflict-resolution rates.
Figure 7: figure
Ablation study for V-RAG, showing that larger $k$ increases the helpful retrieval rate but also raises the harmful rate.
4 Discussion
RadAgents leverages radiologist-inspired workflows, tool-augmented reasoning, and visual retrieval to improve CXR interpretation across three complex benchmarks. However, the system also depends on several external tools, making it sensitive to their errors or distribution shifts. Finally, our evaluation is restricted to offline datasets and automatic metrics; broader human studies and prospective assessments will be needed to understand clinical impact.